OpenStudy (anonymous):

Given this equation of a parabola, does the parabola open upward or downward, and what is its vertex (turning point)? (Hint: First rewrite the equation in standard quadratic form.) y + x2 + 16x = 4

OpenStudy (anonymous):

standart for means put y onn the left and all the rest on the right: y =- x2 - 16x +4 To find the vertex use formula -b/2a to find x-coordinate. -(-16)/2*(-1) = -8 To find y-coordinate substitute in the equation: -64+128 +4 = 68 so vertex is at(-8,68) To see if it opens down or up just look at the sign in front of x2. Since it's negative, opens down

OpenStudy (anonymous):

Thank you so much, I understand it a lot better now :)